# Iterative charge equilibration for fourth-generation high-dimensional neural network potentials

### v1

@@ I Introduction
In spite of many successful applications, second-generation MLPs rely on environment-dependent atomic energies only and therefore do not incorporate any long-range interactions. This gap has been filled by the third generation that extends MLPs by modeling the electrostatic energy and forces through environment-dependent atomic charges Houlding, Liem, and Popelier (2007); Artrith, Morawietz, and Behler (2011); Morawietz, Sharma, and Behler (2012); Unke and Meuwly (2019); Yao et al. (2018); Grisafi and Ceriotti (2019); Gao and Remsing (2022); Zhang et al. (2022). While these potentials offer some improvements in accuracy for some systems Yue et al. (2021), the computational cost is often increased by the need for computing machine learned charges and the explicit calculation of electrostatics, e.g. via an Ewald sum Ewald (1921); Toukmaji and A. B. Board (1996), which can become the bottleneck in large-scale MD simulations. As a comment, it should be noted that the representation of charges and electrostatic multipoles by machine learning has been suggested as early as 2007 Houlding, Liem, and Popelier (2007) to improve the accuracy of the electrostatic energy in classical force fields, i.e., at about the same time as the introduction of second generation MLPs Behler and Parrinello (2007) according to the classification scheme employed here.
@@
In this paper, we introduce the technical details and a benchmark application of the iQEq method, which we have implemented in the open-source software LAMMPS Plimpton (1995), allowing for large-scale MD simulations of periodic systems using the 4G-HDNNP method. This new pair style can be used on multiple cores and is developed as an interface to the n2p2 library Singraber, Behler, and Dellago (2019) for HDNNPs, which to date did not offer the option to run MD simulations based on 4G-HDNNPs. In the next section we provide a concise summary of the 4G-HDNNP method based on conventional dQEq. Then, we describe the iterative iQEq methodology and its implementation in LAMMPS. To illustrate the accuracy and performance of iQEq, we have carried out MD simulations of aqueous FeCl3 solutions with different concentrations and compositions employing a recently reported 4G-HDNNP potential Kocer et al. (2024) for this system. Iron chloride solutions are challenging for MLPs, since the iron ions can adopt two different oxdidation states in solution, Fe2+ and Fe3+. If data for both ions is combined in a single data set, in local second-generation and third-generation MLPs the oxidation state is ill-defined, since chloride ions maybe be located outside the local environments resulting in insufficient information to define the iron oxidation state. As shown in Ref. 38, 4G-HDNNPs are able to overcome this problem and to describe both, aqueous FeCl2 as well as FeCl3 solutions, correctly.
@@ II Methods
The parameters $\sigma_{i}$ determine the widths of the Gaussians representing the atomic charges for the computation of $E_{\mathrm{Coulomb}}$ and $\gamma_{ij}=\sqrt{\sigma_{i}^{2}+\sigma_{j}^{2}}$, whereas $\eta$ is a hyperparameter defining the spatial extention of the real space part in the Ewald sum. To obtain the atomic charges, $E_{\mathrm{QEq}}$ is minimized by solving
@@ II.2 Iterative Charge Equilibration
To improve the computational performance of 4G-HDNNPs, an increased efficiency of the charge equilibration step is central. We address this problem by replacing the cubically scaling dQEq method by an iterative iQEq algorithm, which performs a multidimensional function minimization for calculating the atomic charges and force components. Solving a system of linear equations iteratively via gradient-based optimizations such as the conjugate gradient algorithm has been proposed in the literature before Nakano (1997) and it has been shown that the scaling of the solution can be enhanced significantly. Recently, also Gubler et al. Gubler et al. (2024) discussed this issue and showed that an iterative solution using a rapidly converging conjugate gradient method in combination with a particle mesh solver leads to substantially improved scaling with respect to the number of atoms in the system.
In dQEq, constructing the matrix $\mathbf{A^{\prime}}$ scales quadratically with the number of atoms, while the solution of Eq. 19 exhibits cubic scaling. In the iQEq method, this solution is replaced by a multidimensional function minimization scheme by exploiting that the charge equilibration energy $E_{\mathrm{QEq}}$ is a multidimensional function of the atomic charges
@@
    $\frac{\partial{E_{\mathrm{QEq}}}}{\partial Q_{i}}=\chi_{i}+J_{i}Q_{i}+\frac{2\pi}{V}\sum_{\vec{\mathbf{k}}\neq\mathbf{0}}\frac{e^{\frac{-\eta^{2}|\vec{\mathbf{k}}|^{2}}{2}}}{|\vec{\mathbf{k}}|^{2}}\frac{S(\mathbf{k})\cdot S^{\ast}(\mathbf{k})}{\left|S(\mathbf{k})\right|}+$ (24)
@@
Using the gradient vector, the QEq energy can then be minimized using any gradient-based algorithm. This minimization is an iterative process and the number of iterations depends on the selected gradient algorithm and tolerance values. Two different tolerance parameters need to be selected in the iQEq approach: one is the tolerance for the line minimization and one is the tolerance for the convergence (i.e., gradient tolerance). The former sets the accuracy for the line search process within the minimization algorithm. During each iteration, the algorithm searches along a particular direction defined by the gradient or other optimization criteria to find a step size that minimizes the function in that direction. The latter sets the convergence criterion for the entire minimization process. It ensures that the norm of the gradient falls below a certain threshold, meanining that a local minimum has been reached. In this study, these tolerances are set to 10-2 eV and 10-5 eV/e, respectively, and in general the values to be chosen depend on the desired accuracy and performance as discussed below. It has been observed that the iterative solution significantly improves the scaling of the method with system size, making it attractive for simulations of large systems. The detailed performance comparison between iQEq and dQEq methods is provided in Section III.
@@ III.1 Performance
When measuring the CPU times of the iQEq method it is important to note that a tolerance value has to be chosen defining the accuracy of the iterative charge determination. Moreover, a maximum number of iterations for terminating the iterative minimization needs to be specified. These parameters need to be tested for a given system and depending on the choice of these parameters the absolute timings of the iQEq algorithm are necessarily different. For our particular analysis, we choose 10-5 eV/e as the gradient tolerance and used up to 15 iterations. As discussed in detail below, the required settings depend on the desired accuracy, and here we reach a typical convergence up to the fifth decimal for both atomic charges and forces in units of e and eV/Å, respectively.

### v2

@@ I Introduction
In spite of many successful applications, second-generation MLPs rely on environment-dependent atomic energies only and therefore do not incorporate any long-range interactions. This gap has been filled by the third generation that extends MLPs by modeling the electrostatic energy and forces through environment-dependent atomic charges Houlding, Liem, and Popelier (2007); Artrith, Morawietz, and Behler (2011); Morawietz, Sharma, and Behler (2012); Unke and Meuwly (2019); Yao et al. (2018); Grisafi and Ceriotti (2019); Gao and Remsing (2022); Zhang et al. (2022). While these potentials offer some improvements in accuracy for some systems Yue et al. (2021), the computational cost is often increased by the need for computing machine learned charges and the explicit calculation of electrostatics, e.g. via an Ewald sum Ewald (1921); Toukmaji and A. B. Board (1996), which can become the bottleneck in large-scale molecular dynamics (MD) simulations. As a comment, it should be noted that the representation of charges and electrostatic multipoles by machine learning has been suggested as early as 2007 Houlding, Liem, and Popelier (2007) to improve the accuracy of the electrostatic energy in classical force fields, i.e., at about the same time as the introduction of second generation MLPs Behler and Parrinello (2007) according to the classification scheme employed here.
@@
In this paper, we introduce the technical details and a benchmark application of the iQEq method, which we have implemented in the open-source software LAMMPS Plimpton (1995), allowing for large-scale MD simulations of periodic systems using the 4G-HDNNP method. This new pair style can be used on multiple cores and is developed as an interface to the n2p2 library Singraber, Behler, and Dellago (2019) for HDNNPs, which to date did not offer the option to run MD simulations based on 4G-HDNNPs. In the next section we provide a concise summary of the 4G-HDNNP method based on conventional dQEq. Then, we describe the iterative iQEq methodology and its implementation in LAMMPS. To illustrate the accuracy and performance of iQEq, we have carried out MD simulations of aqueous FeCl3 solutions with different concentrations and compositions employing a recently reported 4G-HDNNP potential Kocer et al. (2024) for this system. Iron chloride solutions are challenging for MLPs, since the iron ions can adopt two different oxidation states in solution, Fe2+ and Fe3+. If data for both ions is combined in a single data set, in local second-generation and third-generation MLPs the oxidation state is ill-defined, since chloride ions may be located outside the local environments resulting in insufficient information to define the iron oxidation state. As shown in Ref. 38, 4G-HDNNPs are able to overcome this problem and to describe both, aqueous FeCl2 as well as FeCl3 solutions, correctly.
@@ II Methods
The parameters $\sigma_{i}$ determine the widths of the Gaussians representing the atomic charges for the computation of $E_{\mathrm{Coulomb}}$ and $\gamma_{ij}=\sqrt{\sigma_{i}^{2}+\sigma_{j}^{2}}$, whereas $\eta$ is a hyperparameter defining the spatial extension of the real space part in the Ewald sum. To obtain the atomic charges, $E_{\mathrm{QEq}}$ is minimized by solving
@@ II.2 Iterative Charge Equilibration
To improve the computational performance of 4G-HDNNPs, an increased efficiency of the charge equilibration step is central. We address this problem by replacing the cubically scaling dQEq method with an iterative iQEq algorithm, which performs a multidimensional function minimization for calculating the atomic charges and force components. Solving a system of linear equations iteratively via gradient-based optimizations such as the conjugate gradient algorithm has been proposed in the literature before Nakano (1997) and it has been shown that the scaling of the solution can be enhanced significantly. Recently, also Gubler et al. Gubler et al. (2024) discussed this issue and showed that an iterative solution using a rapidly converging conjugate gradient method in combination with a particle mesh solver leads to substantially improved scaling with respect to the number of atoms in the system.
In dQEq, constructing the matrix $\mathbf{A^{\prime}}$ scales quadratically with the number of atoms, while the solution of Eq. 19 exhibits cubic scaling. In the iQEq method, this solution is replaced by a multidimensional function minimization scheme exploiting that the charge equilibration energy $E_{\mathrm{QEq}}$ is a multidimensional function of the atomic charges
@@
    $\frac{\partial{E_{\mathrm{QEq}}}}{\partial Q_{i}}=\chi_{i}+J_{i}Q_{i}+\frac{4\pi}{V}\sum_{\vec{\mathbf{k}}\neq\mathbf{0}}\frac{e^{\frac{-\eta^{2}|\vec{\mathbf{k}}|^{2}}{2}}}{|\vec{\mathbf{k}}|^{2}}\mathrm{Re}\left(S(\mathbf{k})e^{i\mathbf{k}\mathbf{r}_{i}}\right)+$ (24)
@@
Using the gradient vector, the QEq energy can then be minimized using any gradient-based algorithm. This minimization is an iterative process and the number of iterations depends on the selected gradient algorithm and tolerance values. Two different tolerance parameters need to be selected in the iQEq approach: one is the tolerance for the line minimization and the other is the tolerance for the convergence (i.e., gradient tolerance). The former sets the accuracy for the line search process within the minimization algorithm. During each iteration, the algorithm searches along a particular direction defined by the gradient or other optimization criteria to find a step size that minimizes the function in that direction. The latter sets the convergence criterion for the entire minimization process. It ensures that the norm of the gradient falls below a certain threshold, meaning that a local minimum has been reached. In this study, these tolerances are set to 10-2 eV and 10-5 eV/e, respectively, and in general the values to be chosen depend on the desired accuracy and performance as discussed below. It has been observed that the iterative solution significantly improves the scaling of the method with system size, making it attractive for simulations of large systems. The detailed performance comparison between iQEq and dQEq methods is provided in Section III.
@@ III.1 Performance
When measuring the CPU times of the iQEq method it is important to note that a tolerance value has to be chosen defining the accuracy of the iterative charge determination. Moreover, a maximum number of iterations for terminating the iterative minimization needs to be specified. These parameters need to be tested for a given system and depending on the choice of these parameters the absolute timings of the iQEq algorithm are necessarily different. For our particular analysis, we choose 10-5 eV/e as the gradient tolerance and used up to 15 iterations. As discussed in detail below, the required settings depend on the desired accuracy, and here we achieve a typical convergence up to the fifth decimal for both atomic charges and forces in units of e and eV/Å, respectively.